Circular No. 6922

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ce IN MCG –4-24-19 F. Patat, European Southern Observatory (ESO); and M. Turatto,

Osservatorio Astronomico di Padova, report that they observed this object (cf. IAUC 6912, 6917) with the ESO 3.6-m telescope (+ EFOSC2; range 330–1000 nm, resolution 1.8 nm) on May 29.01 UT under nonphotometric conditions. A full reduction of the CCD spectrogram shows that this object is a type-II supernova, ∼ 10 days past maximum light. The spectrum is dominated by P-Cyg profiles, superimposed on a relatively blue continuum. The most prominent features are identified as the Balmer lines of hydrogen, Ca II H and K, and a number of Fe II lines. The expansion velocities deduced from the minima of Hα, Hβ, Ca II H and K, and Fe II 516.9nm are 11 200, 9200, 9000, and 7600 km/s, respectively. Estimates of the supernova’s magnitudes are B = 15.9, V = 15.6, and R = 15.3.

XTE J2012+381 F. E. Marshall and T. Strohmayer, Goddard Space Flight Center, on

behalf of the RXTE Science Operations Center, report an improved position for this x-ray transient: α = 20^{h}12^{m}43^{s}, δ = +38^{o}11′.0 (equinox 2000.0); the estimated uncertainty is 1′, based on scans across the source with the Proportional Counter Array on the Rosssi X-ray Timing Explorer on May 27.68 UT. This error circle is consistent with the larger region determined with the RXTE All-Sky Monitor but does not contain the star suggested by Wagner et al. (IAUC 6920). The observed 2–10-keV flux of ∼ 110 mCrab indicates a continued brightening of the source.

M. R. Garcia, J. E. McClintock, P. Berlind, and E. Barton, Harvard-Smithsonian Center for Astrophysics; and P. Callanan, University College, Cork, report: “Spectra and photometry of the possible counterpart to XTE J2012+381 (Wagner et al., IAUC 6920) with the Mt. Hopkins 1.5-m and 1.2m telescopes show that this star is a heavily reddened M giant. Photometry on May 28.312–28.469 UT yields V = 19.7 ± 0.2, R = 16.3 ± 0.1, and I = 12.2 ± 0.1, with no detected variability. The apparent brightening described on IAUC 6920 is most probably an artifact of the extreme red colors and the mismatch in the bandpasses of the LONEOS images (which used an unfiltered CCD) and the DSS images. Neither our photometry nor that on IAUC 6920 provides any evidence that this star is the counterpart to XTE J2012+381.”

1998 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green